They gave away ¼
of what they had to friends and still had 11,462 left. How many did they start with?

Answer by mananth(16949)

let them have x
They gave away 1/4th = x/4
.
x-x/4 =11462
4x-x / 4 = 11462
3x= 45848
x=45848/3
They had about 15,282